# Discrete-Time CRLB-based Power Allocation for CF MIMO-ISAC with Joint Localization and Velocity Sensing

### v2

@@ IV-A3 The algorithm implementation
Input: The positive scalar termination factor $\epsilon_{\mu}$ and $\epsilon_{\rm th}$, small positive constant $\sigma_{\upsilon}$, $\epsilon_{L}$ and $\epsilon_{\upsilon}$, initial penalty factor $\mu_{1}$, scaling factor $\varphi>1$; initial PA parameter ${\bm{\rho}}_{1}$ satisfying ${\bm{1}}^{\rm T}{\bm{\rho}}_{1}=1$, projection operator ${\bm{\Theta}}$, ${\bm{y}}_{1}={\bm{\rho}}_{1}$, and ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})$; a counter $k=0$ and $i=j=1$.
@@
5:  Deflection choice: According to (41), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$;
@@
IV-B Modified Steepest Descent Algorithm
In fact, alternative unconstrained optimization methods can be applied to solve the proposed projection-based penalty problem in (26). In this paper, we also develop a MSD algorithm, adapted from the proposed MCG framework. The main difference between the MSD and MCG algorithms lies in how the search direction is determined. In the MSD approach, the search direction at each iteration is set to the negative gradient of the objective function, i.e.,
@@ IV-B Modified Steepest Descent Algorithm
IV-C Benchmarking Schemes
To facilitate performance comparison, we introduce two benchmark algorithms based on the PP-MCG-ILS and PP-MSD-ILS frameworks, called penalty function and projection-driven normalized SD (PP-NSD) and penalty function and projection-driven normalized CG (PP-NCG), respectively. Here, we employ the standard SD and CG methods proposed in [42], where fixed step sizes are adopted instead of ILS. To mitigate potential instability arising from overly large or small gradient magnitudes under fixed-step updates, we normalize the update directions using their $l_{2}$ norm. Specifically, for the PP-NSD algorithm, the search direction is defined as an $\epsilon_{d}$-regularized normalized gradient, given by
@@ IV-C Benchmarking Schemes
where ${\bm{g}}_{i}={\bm{\Theta}}\nabla L({\bm{\rho}}_{i})$, and $\epsilon_{d}$ is a small positive constant introduced to prevent division by zero. Similarly, for the PP-NCG algorithm, the search direction is given by an $\epsilon_{d}$-regularized normalized conjugate gradient, i.e.,
@@
with $\upsilon_{i}^{\rm b}$ given in (47) and $\upsilon_{0}$ being a predetermined step size. The maximum number of iterations is set to $I$. Due to the non-monotonic penalty function, the minimum may not occur at the final iteration but within a few steps before it. To address this, we introduce an alternative penalty stopping criterion given by $\mu_{j}\min\limits_{i+1-S\leq q\leq i+1}\{\alpha({\bm{\rho}}_{q})\}<\epsilon_{\mu}$ where $S$ represents the window size for searching the minimum penalty value. The detailed implementation steps of the PP-NSD algorithm are omitted here, as it can be easily constructed from the PP-NCG described in Algorithm 2 by removing step 4 and performing the direction update in step 5 using (50).
Algorithm 2 The PP-NCG algorithm
Input: Positive scalar termination factor $\epsilon_{\mu}$ and $\epsilon_{\rm th}$, a large integer $I$ and a small integer $S$, small positive constants $\sigma_{\upsilon}$, $\epsilon_{L}$, $\epsilon_{d}$ and $\epsilon_{\upsilon}$, initial step size $\upsilon_{0}$, initial penalty factor $\mu_{1}$ and its scaling factor $\varphi>1$; initial PA parameter ${\bm{\rho}}_{1}$ satisfying ${\bm{1}}^{\rm T}{\bm{\rho}}_{1}=1$, projection operator ${\bm{\Theta}}$, ${\bm{y}}_{1}={\bm{\rho}}_{1}$, and ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})/(\|{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})\|_{2}+\epsilon_{d})$; a counter $k=0$ and $i=j=1$.
@@
4:  Deflection choice: According to (41), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$;
@@ V Extension to Pure Sensing
In this section, we consider a pure sensing scenario with joint location and velocity estimation. The P-NCG-ILS algorithm is proposed for PA in minimizing the total power under power budget constraints and CRLB constraints.
The constrained power minimization problem is formulated as
@@
$\small\alpha_{\rm s}({\bm{\rho}})=$ ${f}_{\rm s}({\bm{\rho}})+\tilde{f}_{l}({\bm{\rho}})+\tilde{f}_{v}({\bm{\rho}})$
$+\sum_{n=1}^{N}\left(f_{\min,n}({\bm{\rho}})+f_{\max,n}({\bm{\rho}})\right),$ (54)
@@
5:  Deflection choice: According to (41), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$.
@@ VI Algorithm Analysis
The proposed algorithms (e.g., PP-MCG-ILS, PP-MSD-ILS) can be extended and applied to a broad class of nonlinear and non-convex optimization problems. These problems are typically characterized by an affine equality constraint, if present, satisfied through a projection operator, along with nonlinear or non-convex constraints incorporated via penalty functions. The ILS mechanism is used to enhance convergence. For example, the proposed P-NCG-ILS algorithm for sensing-only scenario can be seen as a simple extension of the PP-MCG-ILS algorithm. In the following, we analyze the convergence behavior and computational complexity of the proposed algorithms.
@@ VI-A Convergence Analysis
Theorem 1. Given a nonempty convex set $\Omega$ with the objective function $L({\bm{\rho}}_{i})$ bounded over its interior points, if the objective function is Lipschitz gradient continuous over $\Omega$, the ILS-based iterative algorithm will converge to a stationary point with a zero descent direction ${\bm{d}}_{i}\rightarrow{\bm{0}}$. In contrast, if it is Lipschitz gradient non-continuous, sequences $\{{\bm{\rho}}_{i}\}$ and $\{L({\bm{\rho}}_{i})\}$ may converge to a Lipschitz gradient noncontinuous point with the norm of the Hessian matrix approaching infinity.
@@
Theorem 1 provides the convergence analysis for the proposed iterative algorithms.
We now prove the Lipschitz gradient continuity of the objectivity functions in (26) and (56).
Firstly, all CRLB-involved matrices, e.g., ${\bm{P}}(\rho)$, ${\bm{Y}}(\rho)$, and their inverses are smooth on ${\bm{\rho}}$ over the feasible domain $\Omega\subset\mathbb{R}^{N}$. Moreover, ${\bm{C}}_{L}({\bm{\rho}})$ is uniformly positive definite. It follows that ${\bm{C}}_{L}({\bm{\rho}})$ is continuously differentiable, and so is ${\rm tr}\{{\bm{C}}_{L}({\bm{\rho}})\}$. By the chain rule, we have
@@
where $\nabla f_{l}({\bm{\rho}})=-{\rm tr}\{{\bm{C}}_{L}({\bm{\rho}})\nabla({\bm{P}}-{\bm{V}}{\bm{Y}}^{-1}{\bm{V}}^{\rm T}){\bm{C}}_{L}({\bm{\rho}})\}$. Since the matrix inverse and trace are smooth on the set of positive definite matrices, $\nabla f_{l}({\bm{\rho}})$ is continuously differentiable. Moreover, all matrices involved (including ${\bm{C}}_{L}({\bm{\rho}})$) have bounded spectral norm, so the operator norm of the Hessian of $f_{l}({\bm{\rho}})$ is also bounded. Consequently, $\nabla\tilde{f}_{l}({\bm{\rho}})$ is Lipschitz continuous. A similar argument applies to $\nabla\tilde{f}_{v}({\bm{\rho}})$. On the other hand, $f_{\min,n}({\bm{\rho}})$, $f_{\max,n}({\bm{\rho}})$ and ${f}_{\rm s}({\bm{\rho}})$ are piecewise quadratic, and their gradients are Lipschitz continuous. Hence, $\nabla_{\bm{\rho}}\alpha({\bm{\rho}})$ and $\nabla_{\bm{\rho}}\alpha_{\rm s}({\bm{\rho}})$ are Lipschitz continuous over $\Omega$, and the same applies to $\nabla_{\bm{\rho}}L({\bm{\rho}})$ and $\nabla_{\bm{\rho}}L_{\rm s}({\bm{\rho}})$.
Additionally, since ${\bm{\Theta}}$ is a constant matrix with $\|{\bm{\Theta}}\|_{2}=1$, and $\nabla_{\bm{\rho}}f({\bm{\rho}})$ is Lipschitz continuous, the composition $\nabla\mathcal{L}({\bm{\varpi}})={\bm{\Theta}}\nabla L({\bm{\rho}})$ is also Lipschitz continuous. In other words, there exists a constant $C>0$ such that
@@ VI-B Complexity Analysis
We observe that the increased computations of the PP-MCG-ILS algorithm, compared to the PP-MSD-ILS algorithm, primarily stems from the computation of $\varsigma_{i}^{\rm b}$ and $\varsigma_{i}^{\rm HS}$ at each iteration, which involves approximately $4N$ multiplications and 2 divisions. In both algorithms, the dominant computational burden lies in evaluating the objective function during the ILS and computing the gradients for direction updates, which are inherently tied to the CRLB expressions. Therefore, a more meaningful comparison of the overall computational cost focuses on the number of ILS steps performed per iteration. This number is influenced by the step size threshold, given by the right-hand side of (C), the initial step size $\upsilon_{0,i}$ in (48), and the scaling factor. At a given iterate ${\bm{\rho}}_{i}$, a larger search direction norm $\|{\bm{d}}_{i}\|$ generally leads to a smaller $\upsilon_{0,i}$ and a reduced step size threshold. As a consequence, the number of ILS steps per iteration is expected to remain comparable between the PP-MCG-ILS and PP-MSD-ILS algorithms. Accordingly, the primary metrics for evaluating the relative performance of these two algorithms are their convergence rate and the quality of their steady-state solutions.
We now qualitatively compare the computational complexity between algorithms employing ILS and those without ILS. As previously discussed, the number of ILS steps significantly impacts the number of evaluations of the objective function. Thus, the computational complexity for step size selection of the benchmark algorithms can be considerably lower at each iteration compared to the PP-MCG-ILS and PP-MSD-ILS algorithms. However, all four algorithms have similar complexity in direction update due to the inherently intensive gradient calculations involved. Moreover, the overall computational complexity heavily depends on the convergence rates, or equivalently, on the total number of iterations required for convergence. As will be demonstrated in the subsequent simulation results, the proposed PP-MCG-ILS algorithm exhibits a notably higher convergence rate compared to its counterparts, with or without ILS. This superior convergence performance significantly reduces the overall computational effort, particularly in the sense that it benefits demanding gradient computations.
@@ VII Simulation Results
In this section, we conduct numerical simulations to evaluate the proposed PP-MCG-ILS, PP-MSD-ILS and P-NCG-ILS algorithms for PA optimization. All APs are assumed to be stationary. We employ the orthogonal chirp division multiplexing (OCDM) [44] waveform with a single Gaussian pulse shaping. The carrier frequency is $3$GHz. The lowpass equivalent is formulated by
@@ VII-A CRLB Validation
We first verify the CRLB in comparison with the MLE results. To maintain consistency with existing works [28, 38, 29], we use the signal energy-to-noise ratio (SENR), defined as $\tilde{\delta}_{\rm cn}\triangleq P/\sigma_{\rm cn}^{2}$, instead of the SCNR $\delta_{\rm cn}$ in evaluating the CRLBs. Due to the high complexity of four-dimensional MLE, we consider a simple $4\times 3$ nonsymmetric multi-static radar configuration given in Fig. 2. Without loss of generality, we consider the OCDM waveform with $M=16$, $T=10^{-2}$ s and a sampling rate of $f_{s}=1$ kHz.
Fig. 3 compares the mean squared error (MSE) of the MLE with the derived CRLB. It is observed that the derived CRLBs closely match the MSEs for SENR values above $-20$ dB. The small discrepancies in the low-SENR regime are attributed to the suboptimal behavior of the likelihood function under noise-dominated conditions. Overall, these results validate the effectiveness and accuracy of the proposed CRLBs.

### v3

@@ IV-C3 The algorithm implementation
Input: The positive scalar termination factor $\epsilon_{\mu}$ and $\epsilon_{\rm th}$, small positive constant $\sigma_{\upsilon}$, $\epsilon_{L}$ and $\epsilon_{\upsilon}$, initial penalty factor $\mu_{1}$, scaling factor $\varphi>1$; initial PA parameter ${\bm{\rho}}_{0}$ satisfying ${\bm{1}}^{\rm T}{\bm{\rho}}_{0}=1$, projection operator $\bm{\Theta}$}, ${\bm{y}}_{1}={\bm{\rho}}_{1}$, and ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})$; a counter $k=0$ and $i=j=1$.
@@
5:  Deflection choice: According to (41), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=Q$, let $\varsigma_{i}=0$ and $k=0$;
@@
IV-D Modified Steepest Descent Algorithm
In fact, alternative unconstrained optimization methods can be applied to solve the proposed projection-based penalty problem in (26). In this paper, we also develop an MSD algorithm, adapted from the proposed MCG framework. The main difference between the MSD and MCG algorithms lies in how the search direction is determined. In the MSD approach, the search direction is set to the projected negative gradient of the penalty-based objective function, i.e.,
@@ IV-D Modified Steepest Descent Algorithm
IV-E Benchmarking Schemes
Here, we employ the standard SD and CG methods proposed in [37] as benchmarks, where fixed step sizes are adopted instead of ILS. To facilitate performance comparison, these two benchmark algorithms should be adapted to our problem below, called penalty function and projection-driven normalized SD (PP-NSD) and penalty function and projection-driven normalized CG (PP-NCG), respectively.
To mitigate potential instability arising from overly large or small gradient magnitudes under fixed-step updates, we normalize the update directions of CG or SD algorithms using their $l_{2}$ norm. Specifically, for the PP-NSD algorithm, the search direction is expressed by an $\epsilon_{d}$-regularized normalized gradient
@@ IV-E Benchmarking Schemes
where ${\bm{g}}_{i}={\bm{\Theta}}\nabla L({\bm{\rho}}_{i})$, and $\epsilon_{d}$ is a small positive constant to prevent division by zero. Similarly, for the PP-NCG algorithm, the search direction is given by an $\epsilon_{d}$-regularized normalized CG, i.e.,
@@
with $\upsilon_{i}^{\rm b}$ given in (47) and $\upsilon_{0}$ being a predetermined step size. Due to the uncertain convergence of the benchmarks, their maximum number of iterations is set to $I$. Moreover, due to the non-monotonic penalty function, the minimum may not occur at the final iteration but within a few steps before it. To address this, we introduce an alternative penalty stopping criterion given by $\mu_{j}\min\limits_{i+1-S\leq q\leq i+1}\{\alpha({\bm{\rho}}_{q})\}<\epsilon_{\mu}$ where $S$ represents the window size for searching the minimum penalty value. The detailed implementation steps of the PP-NSD algorithm are omitted here, as it can be easily constructed from the PP-NCG described in Algorithm 2 by removing step 4 and performing the direction update in step 5 using (50).
Algorithm 2 The $P$-NCG algorithm
Input: Positive scalar termination factor $\epsilon_{\mu}$ and $\epsilon_{\rm th}$, a large integer $I$ and a small integer $S$, small positive constants $\sigma_{\upsilon}$, $\epsilon_{L}$, $\epsilon_{d}$ and $\epsilon_{\upsilon}$, initial step size $\upsilon_{0}$, initial penalty factor $\mu_{1}$ and its scaling factor $\varphi>1$; initial PA parameter ${\bm{\rho}}_{0}$ satisfying ${\bm{1}}^{\rm T}{\bm{\rho}}_{0}=1$, projection operator $\bm{\Theta}$}, ${\bm{y}}_{1}={\bm{\rho}}_{1}$, and ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})/(\|{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})\|_{2}+\epsilon_{d})$; a counter $k=0$ and $i=j=1$.
@@
4:  Deflection choice: According to (41), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=Q$, let $\varsigma_{i}=0$ and $k=0$;
@@ V Extension to Pure Sensing
In this section, we consider a pure sensing scenario with joint location and velocity estimation. The P-NCG-ILS algorithm is proposed for PA optimization in minimizing the total power under power budget constraints and CRLB constraints.
The constrained power minimization problem is given by
@@
$\small\alpha_{\rm s}({\bm{\rho}})=\$${f}_{\rm s}({\bm{\rho}})+\tilde{f}_{l}({\bm{\rho}})+\tilde{f}_{v}({\bm{\rho}})$
$+\sum\nolimits_{n=1}^{N}\left(f_{\min,n}({\bm{\rho}})+f_{\max,n}({\bm{\rho}})\right),$ (54)
@@
5:  Deflection choice: According to (41), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=Q$, let $\varsigma_{i}=0$ and $k=0$.
@@ VI Algorithm Analysis
The proposed algorithms (e.g., PP-MCG-ILS and PP-MSD-ILS) can be readily extended to a broad class of nonlinear and nonconvex optimization problems. Such problems typically involve affine equality constraints, which, when present, are enforced via projection operators, while nonlinear or nonconvex constraints are handled through penalty functions. The ILS mechanism is further employed to improve convergence stability. As an example, the proposed P-NCG-ILS algorithm for the sensing-only scenario can be regarded as a direct extension of the PP-MCG-ILS algorithm.
In the following, we analyze the convergence behavior and computational complexity of the proposed algorithms.
@@ VI-A Convergence Analysis
Theorem 1. Consider the CRLB-based penalty objective $L(\bm{\rho})$ defined in (24), over a nonempty and bounded convex feasible set $\Omega\subset\mathbb{R}^{n}$. If $L$ is continuously differentiable and admits a Lipschitz continuous gradient on $\Omega$, the proposed ILS-based schemes (PP-MCG-ILS and PP-MSD-ILS) generate a sequence $\{\bm{\rho}_{i}\}$ whose limit points are first-order stationary with a zero descent direction ${\bm{d}}_{i}\rightarrow{\bm{0}}$.
Theorem 1 is included to guarantee the existence of a well-defined and bounded step size for the proposed ILS-based PA algorithms.
@@
We now establish the Lipschitz continuity of the gradients of the objective functions in (26) and (56) by showing that their second-order derivatives are uniformly bounded over the feasible set.
Firstly, since $\Omega\subset\mathbb{R}^{N}$ is compact and all CRLB-related matrices, e.g., ${\bm{P}}(\rho)$, ${\bm{Y}}(\rho)$, and their inverses depend smoothly on $\bm{\rho}$}, their spectral norms and those of their derivatives are uniformly bounded over $\Omega$. As a result, the operator norm of the Hessian of $f_{l}({\bm{\rho}})$ is uniformly bounded.
Moreover, we know that ${\bm{C}}_{L}({\bm{\rho}})$ is uniformly positive definite. It follows that ${\bm{C}}_{L}({\bm{\rho}})$ is continuously differentiable, and so is ${\rm tr}\{{\bm{C}}_{L}({\bm{\rho}})\}$. By the chain rule, we have
@@
where $\nabla f_{l}({\bm{\rho}})=-{\rm tr}\{{\bm{C}}_{L}({\bm{\rho}})\nabla({\bm{P}}-{\bm{V}}{\bm{Y}}^{-1}{\bm{V}}^{\rm T}){\bm{C}}_{L}({\bm{\rho}})\}$. Since the matrix inverse and trace are smooth on the set of positive definite matrices, $\nabla f_{l}({\bm{\rho}})$ is continuously differentiable. Together with the uniform boundedness of the Hessian $\nabla^{2}f_{l}({\bm{\rho}})$ over $\Omega$, this implies that $\nabla f_{l}({\bm{\rho}})$ is Lipschitz continuous over $\Omega$. Consequently, $\nabla\tilde{f}_{l}({\bm{\rho}})$ is Lipschitz continuous. A similar argument applies to $\nabla\tilde{f}_{v}({\bm{\rho}})$.
Furthermore, $f_{\min,n}({\bm{\rho}})$, $f_{\max,n}({\bm{\rho}})$ and ${f}_{\rm s}({\bm{\rho}})$ are piecewise quadratic with finitely many breakpoints, and their gradients are globally Lipschitz continuous over $\Omega$. Hence, $\nabla_{\bm{\rho}}\alpha({\bm{\rho}})$ and $\nabla_{\bm{\rho}}\alpha_{\rm s}({\bm{\rho}})$ are Lipschitz continuous over $\Omega$, and the same applies to $\nabla_{\bm{\rho}}L({\bm{\rho}})$ and $\nabla_{\bm{\rho}}L_{\rm s}({\bm{\rho}})$.
Finally, since $\bm{\Theta}$} is a constant matrix with $\|{\bm{\Theta}}\|_{2}=1$, and $\nabla_{\bm{\rho}}f({\bm{\rho}})$ is Lipschitz continuous, the composition $\nabla\mathcal{L}({\bm{\varpi}})={\bm{\Theta}}\nabla L({\bm{\rho}})$ is also Lipschitz continuous. In other words, there exists a constant $C>0$, independent of $\bm{\varpi}$}, such that
@@ VI-B Complexity Analysis
We observe that the increased computations of the PP-MCG-ILS algorithm, compared to the PP-MSD-ILS algorithm, primarily stems from the computation of $\varsigma_{i}^{\rm b}$ and $\varsigma_{i}^{\rm HS}$ at each iteration, which involves approximately $4N$ multiplications and 2 divisions. In both algorithms, the dominant computational burden lies in evaluating the objective function during the ILS and computing the gradients for direction updates, which are inherently tied to the CRLB expressions. Therefore, a more meaningful comparison of the overall computational cost focuses on the number of ILS steps performed per iteration. This number is influenced by the step size threshold, given by the right-hand side of (C), the initial step size $\upsilon_{i,0}$ in (48), and the scaling factor. At a given iterate ${\bm{\rho}}_{i}$, a larger search direction norm $\|{\bm{d}}_{i}\|$ generally leads to a smaller $\upsilon_{i,0}$ and a reduced step size threshold. As a consequence, the number of ILS steps per iteration is expected to remain comparable between the PP-MCG-ILS and PP-MSD-ILS algorithms. Therefore, their relative performance is primarily characterized by their convergence rates and the quality of the converged solutions.
We next compare algorithms with and without ILS. As discussed above, the number of ILS steps directly affects the number of objective-function evaluations. Accordingly, benchmark algorithms without ILS incur a lower step size selection cost per iteration compared with the PP-MCG-ILS and PP-MSD-ILS algorithms. However, all four algorithms exhibit similar computational complexity in the direction update, since gradient evaluations dominate this stage.
Finally, the overall computational complexity is strongly influenced by the convergence rate, or equivalently, by the total number of iterations required to reach convergence. As demonstrated in the simulation results, the proposed PP-MCG-ILS algorithm converges significantly faster than the benchmark methods, both with and without ILS. This faster convergence substantially reduces the overall computational effort, particularly by alleviating the repeated evaluation of computationally intensive gradients.
@@ VII Simulation Results
In this section, we conduct numerical simulations to evaluate the proposed PP-MCG-ILS, PP-MSD-ILS and P-NCG-ILS algorithms for PA optimization. All APs are assumed to be stationary. We employ the orthogonal chirp division multiplexing (OCDM) [31] waveform with a single Gaussian pulse shaping. The carrier frequency is 33GHz. The lowpass equivalent is formulated by
@@ VII-A CRLB Validation
To verify the approximate CRLB, we first compare it with the mean squared error (MSE) of MLE results across the considered SCNR range. To maintain consistency with existing works [16, 17, 3], we use the signal energy-to-noise ratio (SENR), defined as $\tilde{\delta}_{\rm cn}\triangleq P/\sigma_{\rm cn}^{2}$, instead of the SCNR $\delta_{\rm cn}$ in evaluating the CRLBs. Due to the high complexity of four-dimensional MLE, we consider a simple $4\times 3$ nonsymmetric multi-static radar configuration given in Fig. 2. Without loss of generality, we consider the OCDM waveform with $M=16$, $T=10^{-2}$ s and a sampling rate of $f_{s}=1$ kHz.
Fig. 3 compares the MSE of the MLE with the derived CRLB. It is observed that the derived CRLBs closely match the MSEs for SENR values above 20-20 dB. The small discrepancies in the low-SENR regime are attributed to the suboptimal behavior of the likelihood function under noise-dominated conditions. Overall, these results validate the effectiveness and accuracy of the proposed CRLBs.